# ACPBench: Reasoning about Action, Change, and Planning

### v1

@@ Abstract
There is an increasing body of work using Large Language Models (LLMs) as agents for orchestrating workflows and making decisions in domains that require planning and multi-step reasoning. As a result, it is imperative to evaluate LLMs on core skills required for planning. In this work, we present ACPBench, a benchmark for evaluating the reasoning tasks in the field of planning. The benchmark consists of $7$ reasoning tasks over $13$ planning domains. The collection is constructed from planning domains described in a formal language. This allows us to synthesize problems with provably correct solutions across many tasks and domains. Further, it allows us the luxury of scale without additional human effort, i.e., many additional problems can be created automatically. Our extensive evaluation of $22$ open-sourced and frontier LLMs highlight the significant gap in the reasoning capability of the LLMs. The average accuracy of one of the best-performing frontier LLMs — GPT-4o on these tasks can fall as low as $52.50\%$ ACPBench collection is available at the following link: https://ibm.github.io/ACPBench
@@ 1 Introduction
Figure 1: Performance of few state-of-the-art LLMs over different tasks in ACPBench. While the largest LLMs achieve more than $80\%$ accuracy on few tasks, the variance in performance across tasks and across LLMs is still big. This signifies the long way to go before they can be reliably used in practical scenarios.
@@
We evaluate performance of $22$ state-of-the-art language models including open-sourced Phi-3 128K (Abdin et al. 2024), Mixtral 8x22B (MistralAI 2024), LLAMA-3 70B (Dubey et al. 2024), and a closed source GPT-4o (OpenAI 2024) on the ACPBench. We found that, even with Chain-of-Thought prompting (COT)  (Wei et al. 2022) and 2-shot examples, GPT-4o was only able to achieve $78.40\%$ accuracy on MCQ questions in the ACPBench; which reduced to $52.50\%$ for the most difficult task. To understand whether the smaller language models can improve their performance on these tasks, we fine-tune a language model on these tasks. The fine-tuning resulted in substantial improvements in performance across tasks and even demonstrated the ability to generalize to previously unseen domains.
@@
  • We evaluate $22$ state-of-the-art language models of different sizes on ACPBench; including closed source GPT-4o and GPT-4o Mini.
@@ 2 Related Work and Background
Recognizing the importance of evaluating reasoning and planning ability of LLMs, various benchmarks have been proposed (Liu et al. 2023; Ma et al. 2024). Most relevant to our work are the benchmarks that are generated from PDDL tasks. He et al. (2023) proposed a natural language based question answering style dataset to evaluate LLMs on 4 tasks of projection, execution, planning, and goal recognition. PlanBench (Valmeekam et al. 2023b) is a benchmark suite with $8$ planning tasks including plan generation, reasoning about plan execution, plan verification etc. Both these benchmarks focus majorly on the BlocksWorld domain, and employ a template-based approach to generate natural language text. In contrast, AutoPlanBench (Stein et al. 2024) proposes to leverage LLMs to generate the natural language template. Specifically, they propose to prompt LLM for natural language template for each predicate and action. By reducing the human effort required for template generation, they were able to scale up and generate a dataset of $12$ domains.
[…]
3 ACPBench
@@ 4. Action Reachability (AReach)
The action reachability task is closely related to the atom reachability. If an action model is available, then action reachability is equivalent to the atom reachability over the preconditions of the action. Therefore, this task requires an additional reasoning step about action preconditions. Similarly to the atom reachability task, the positive examples are generated from action rollouts, while the negative examples are generated by collecting actions with preconditions including unreachable atoms according to two of the three cases mentioned above delete-relaxed reachability and mutexes. The third case, unreachable static facts, was not used as often creates non-sensible actions board car l0 at location c1. Instead, we added incorrect action templates for each action, like “board the car c1 at location l0 into the airplane” or “drive from location l0 to location l1”. Here as well, the chains of thoughts are created in a similar manner, and the MCQ is captured based on the positive and negative options lists.
@@ 7. Landmarks (Land)
While checking whether a fact is a landmark is PSPACE-hard (Porteous, Sebastia, and Hoffmann 2001), there are several methods that can find a subset of landmarks (Keyder, Richter, and Helmert 2010; Hoffmann, Porteous, and Sebastia 2004; Richter, Helmert, and Westphal 2008; Zhu and Givan 2003). We use the so-called RHW method (Richter, Helmert, and Westphal 2008). Further, negative evidence can be obtained from a collection of plans - a fact that does not appear on all of these plans is not a landmark. We sample from positive and negative examples obtained that way and construct two boolean questions and one MCQ. Here as well, the chains of thoughts generated capture the described logic.
@@ 4 Experiments
4.1 Evaluation of pre-trained models
@@
We first analyse how existing pre-trained models perform on ACPBench. Table 2 presents the accuracy of all the language models on the $7$ ACPBench tasks. These results are mean over $5$ runs for all models; except GPT family models and LLAMA-3.1 405B (Dubey et al. 2024), which were run once due to resource constraints. All LLMs were either accessed using API or hosted locally using hugging face transformer library on machines with 2 $A100$ $80$GB GPU. Note that accuracy of $50.00$ on boolean questions indicates that the performance of the model is as good as a random guess. As all the MCQs in the datasets have $4$ options, accuracy less than $25.00$ indicates that the performance is worse than random guess. LLMs were evaluated on these questions with two in-context examples and Chain-of-Thought (COT) prompting. An example prompt for the Bool applicability question is shown in Figure 3. Note that we use 2-shot but only show 1-shot in the example due to space constraint.
@@
Figure 4 displays a domain-wise analysis of the performance of LLMs on multi-choice questions. This analysis showcases the top $8$ performing models [^3] All supplementary information on the remaining models and boolean questions are relegated to the Appendix to maintain clarity.. The average performance of these top-$8$ models is shown in Fig. 4 as the dotted line in black. This indicates that across models no specific domain seems too easy. However, Rovers, FloorTile, Blocksworld, Alfworld and Satellite domains pose the greatest challenges to LLMs, in that particular order.
@@ 4.2 Fine-tuning
Upon finetuning, the average accuracy of the model improves from $51.43\%$ to $95.71\%$ on boolean questions and from $19.18\%$ to $94.29\%$ on multi-choice questions. Further, Table 4 presents the performance on the remaining $5$ unseen domains. It is remarkable to observe such a significant improvement even on unseen domains; sometimes surpassing the GPT-4o performance. This indicates that finetuning a model, even on a separate domain, improves performance on these tasks. The right-most column in Tables 3 and 4 presents the performance of the best on that task LLM with COT 2-shot prompting. As can be seen; Granite Finetuned model outperforms the best of all models for most of the tasks in the training domains. Even in testing domains, the accuracy difference is significantly reduced upon finetuning.
@@ 4.3 Ablations
Task Base IO Base COT 2shot Finetuned IO Best
@@
Table 3: Comparison of the Granite-code Base 8B model and the finetuned model on $8$ training domains of ACPBench. We present accuracy values for the Base model with Input-Output prompts (w/ IO) as well as with chain-of-thought prompt with two in-context examples (COT 2-shot). The values enclosed in parentheses represent the improvement over the base model w/ IO prompts. The right-most column presents the performance of the best LLM with COT 2-shot on training domain. Best results are in bold.
Task Base IO Base COT 2shot Finetuned IO Best
@@ 5 Discussion and Future Work
In this work, we introduce ACPBench—a collection of datasets to evaluate the ability of LLMs to reason about action, change and planning. By evaluating $22$ state-of-the-art LLMs of varying size, we find these models to underperform, even the largest ones, especially on tasks like plan validation and action reachability. On the other hand, we show that finetuning a small language model, Granite 8B, can improve its reasoning ability to bring it on par with the best performing models. Further, we observe that the fine-tuned model exhibits remarkable generalization ability to unseen domains in ACPBench as well as to a different task in PlanBench.
@@ A.2 Pretrained LLMs
Our paper presents domain-wise performance of few selected models in Fig 4. We only presented results for the MCQ questions due to space constraints. Performance of all the $22$ pretrained LLMs for both boolean and multi-choice questions is presented in Table 6. Table 2 presents accuracy values with 2-shot COT prompting. We also attempted 0-shot Input Output prompt, the trends were similar. These results aggregated over $5$ runs are presented in Table 7.
@@ A.3 Finetuned LLM
In Tables 3 and 4, we compare te finetuned model against the best performing model (last column). These values are obtained by looking at the aggregated performance on train and test domains respectively. We present performance of all the models on train and test domains in Tables 10 and  11.
@@ A.4 Ablation: Prompt Style
In our paper, we present an ablation to compare prompting styles. For that analysis; we presented results on test domains for multi-choice questions in Fig 5. Here, in Fig. 7, we present performance on boolean questions.
@@
Figure 8: Comparison of different prompt styles on LLAMA 3.1 405B for Boolean tasks & MCQ task in all ACPBench domains.
@@
[ caption = Example questions for $7$ ACPBench tasks. For each question, an LLM was provided with context and then the question. The context contains natural language descriptions., label = tab:example_tasks, ] colspec = c—c—X—X, width = Task Context Boolean Questions MCQ Question
@@
Table 6: Accuracy values for $22$Pretrained LLMs on Boolean and MCQ questions in ACPBench; segregated per domain.
@@
Table 7: Accuracy of $22$ LLMs on $7$ ACPBench tasks (boolean as well as multi-choice questions) with Zeroshot IO prompt. The best results are $\mathbf{boldfaced}$, second best are $\underline{underlined}$, and the best among the small, open-sourced models are highlighted with *. The right-most column is mean across tasks.

### v2

@@ Abstract
There is an increasing body of work using Large Language Models (LLMs) as agents for orchestrating workflows and making decisions in domains that require planning and multi-step reasoning. As a result, it is imperative to evaluate LLMs on core skills required for planning. In this work, we present ACPBench, a benchmark for evaluating the reasoning tasks in the field of planning. The benchmark consists of $7$ reasoning tasks over $13$ planning domains. The collection is constructed from planning domains described in a formal language. This allows us to synthesize problems with provably correct solutions across many tasks and domains. Further, it allows us the luxury of scale without additional human effort, i.e., many additional problems can be created automatically. Our extensive evaluation of $22$ LLMs and OpenAI o1 reasoning models highlight the significant gap in the reasoning capability of the LLMs. Our findings with OpenAI o1, a multi-turn reasoning model, reveal significant gains in performance on multiple-choice questions, yet surprisingly, no notable progress is made on boolean questions.
ACPBench collection is available at the following link: https://ibm.github.io/ACPBench
@@ 1 Introduction
Figure 1: Performance of few state-of-the-art LLMs and OpenAI o1 reasoning models over different tasks in ACPBench. While the largest LLMs achieve more than $80\%$ accuracy on few tasks, the variance in performance across tasks and across LLMs is still big. This signifies the long way to go before they can be reliably used in practical scenarios.
@@
We evaluate performance of OpenAI o1 reasoning model and $22$ state-of-the-art language models (including open-sourced Phi-3 128K (Abdin et al. 2024), Mixtral 8x22B (MistralAI 2024), LLAMA-3 70B (Dubey et al. 2024), and a closed source GPT-4o (OpenAI 2024a)) on the ACPBench. We found that, with Chain-of-Thought prompting (COT)  (Wei et al. 2022) and 2-shot examples, GPT-4o was only able to achieve $78.40\%$ accuracy on MCQ questions in the ACPBench; with lowest accuracy of $52.50\%$ for the most difficult (validation) task. Similarly, OpenAI o1 preview achieves accuracy of $87.31\%$ on average for the MCQ questions, with lowest accuracy of $63.08\%$ for the most difficult task. Figure 1 shows the overall performance of few selected models on all $7$ tasks of ACPBench. To understand whether the smaller language models can improve their performance on these tasks, we fine-tune a language model on these tasks. The fine-tuning resulted in substantial improvements in performance across tasks and even demonstrated the ability to generalize to previously unseen domains.
@@
  • We evaluate OpenAI o1 reasoning models and $22$ state-of-the-art language models of different sizes on ACPBench.
@@ 2 Related Work and Background
Recognizing the importance of evaluating reasoning and planning ability of LLMs, various benchmarks have been proposed (Liu et al. 2023; Ma et al. 2024). Most relevant to our work are the benchmarks that are generated from PDDL tasks. He et al. (2023) proposed a natural language based question answering style dataset to evaluate LLMs on 4 tasks of projection, execution, planning, and goal recognition. PlanBench (Valmeekam et al. 2023b) is a benchmark suite with $8$ planning tasks including plan generation, reasoning about plan execution, and plan verification. Both these benchmarks focus on a limited number of planning domains (mainly the BlocksWorld domain), employing a template-based approach to generate natural language text. In contrast, AutoPlanBench (Stein et al. 2024) proposes to leverage LLMs to generate the natural language template. Specifically, they prompt an LLM for natural language template per predicate and per action. By reducing the human effort required for template generation, they were able to scale up the dataset to $12$ domains. However, they limit their focus to a single task - plan generation.
@@
In parallel, Handa et al. (2024) proposed ActionReasoningBench, featuring six tasks: Fluent Tracking, State Tracking, Action Executability, Effects of Actions, Numerical RAC, and Composite Questions. Although there is some overlap between the tasks in ActionReasoningBench and ACPBench (for example, the Effects of Actions task overlaps with our Progression task), the majority of the tasks we propose are not covered by ActionReasoningBench: Reachability, Action Reachability, Validation, Justification, Landmarks. Similarly, the following ActionReasoningBench tasks are not covered in ACPBench: State Tracking, and Numerical RAC.
[…]
3 ACPBench
3.1 Domains
3.2 ACPBench Tasks
@@ 4. Action Reachability (AReach)
The action reachability task is closely related to the atom reachability. If an action model is available, then action reachability is equivalent to the atom reachability over the preconditions of the action. Therefore, this task requires an additional reasoning step about action preconditions. Similarly to the atom reachability task, the positive examples are generated from action rollouts, while the negative examples are generated by collecting actions with preconditions including unreachable atoms according to two of the three cases mentioned above delete-relaxed reachability and mutexes. The third case, unreachable static facts, was not used as often creates non-sensible actions board car l0 at location c1. Instead, we added incorrect action templates for each action, like “board the car c1 at location l0 into the airplane” or “drive from location l0 to location l1”. Here as well, the chain of thoughts are created in a similar manner, and the MCQ is captured based on the positive and negative options lists.
@@ 7. Landmarks (Land)
While checking whether a fact is a landmark is PSPACE-hard (Porteous, Sebastia, and Hoffmann 2001), there are several methods that can find a subset of landmarks (Keyder, Richter, and Helmert 2010; Hoffmann, Porteous, and Sebastia 2004; Richter, Helmert, and Westphal 2008; Zhu and Givan 2003). We use the so-called RHW method (Richter, Helmert, and Westphal 2008). Further, negative evidence can be obtained from a collection of plans - a fact that does not appear on all of these plans is not a landmark. We sample from positive and negative examples obtained that way and construct two boolean questions and one MCQ. Here as well, the chain of thoughts generated capture the described logic.
@@ 4 Experiments
4.1 Evaluation of pre-trained language models
@@
We first analyse how existing pre-trained language models perform on ACPBench. Table 2 presents the accuracy of all the language models on the $7$ ACPBench tasks. These results are mean over $5$ runs for all models; except GPT family models and LLAMA-3.1 405B (Dubey et al. 2024), which were run once due to resource constraints. All LLMs were either accessed using API or hosted locally using hugging face transformer library on machines with 2 $A100$ $80$GB GPU. Note that accuracy of $50.00$ on boolean questions indicates that the performance of the model is as good as a random guess. As all the MCQs in the datasets have $4$ options, accuracy less than $25.00$ indicates that the performance is worse than random guess. To investigate the out-of-the-box performance, we restrict the evaluation to single turn Chain-of-Thought (COT) prompting with two in-context examples. An example prompt for the Bool applicability question is shown in Figure 3.
@@
Figure 4 displays a domain-wise analysis of the performance of LLMs on multi-choice questions. This analysis showcases the top $8$ performing models [^3] All supplementary information on the remaining models and boolean questions are relegated to the Appendix to maintain clarity.. The average performance of these top-$8$ models is shown in Figure 4 as the dotted line in black. This indicates that across models no specific domain seems too easy. However, Rovers, FloorTile, Blocksworld, Alfworld and Satellite domains pose the greatest challenges to LLMs, in that particular order.
@@ 4.2 Fine-tuning
Upon finetuning, the average accuracy of the model improves from $51.43\%$ to $95.71\%$ on boolean questions and from $19.18\%$ to $94.29\%$ on multi-choice questions. Further, Table 4 presents the performance on the remaining $5$ unseen domains. It is remarkable to observe such a significant improvement even on unseen domains; sometimes surpassing the GPT-4o performance. This indicates that finetuning a model, even on a separate domain, improves performance on these tasks. The right-most column in Tables 3 and 4 presents the performance of the best on that task LLM with COT 2-shots prompting. As can be seen; Granite Finetuned model outperforms the best of all models for most of the tasks in the training domains. Even in testing domains, the accuracy difference is significantly reduced upon finetuning.
@@ 4.3 Ablations
Task Base IO Base COT 2-shot Finetuned IO Best
@@
Table 3: Comparison of the Granite-code Base 8B model and the finetuned model on $8$ training domains of ACPBench. We present accuracy values for the Base model with Input-Output prompts (IO) as well as with Chain-of-Thought prompt with two in-context examples (COT 2-shot). The values enclosed in parentheses represent the improvement over the base model w/ IO prompts. The right-most column presents the performance of the best LLM with COT 2-shot on training domain. Best results are in bold.
Task Base IO Base COT 2-shot Finetuned IO Best
@@
4.4 Reasoning Model: OpenAI o1
(a) Bool
(b) MCQ
Figure 7: Comparing OpenAI o1 models with the best LLM. Positive difference shows OpenAI o1 model performing better than the best of the LLMs. Negative difference is when OpenAI o1 model lags behind the best LLM.
Bool MCQ Bool MCQ Bool MCQ Bool MCQ Bool MCQ Bool MCQ Bool MCQ Bool MCQ
2-shot Chain-of-Thought prompt
LLAMA 405B $95.38$ $86.92$ $93.08$ $93.85$ $59.23$ $80.77$ $77.23$ $62.92$ $65.00$ $65.00$ $90.00$ $86.92$ $83.08$ $65.38$ $80.43$ $77.39$
GPT-4o Mini $90.77$ $73.85$ $95.38$ $79.23$ $80.77$ $39.23$ $67.69$ $46.15$ $54.17$ $21.67$ $77.69$ $70.00$ $76.92$ $67.69$ $77.63$ $56.83$
GPT-4o $96.92$ $89.23$ $94.62$ $90.00$ $79.23$ $76.92$ $61.54$ $53.85$ $57.50$ $52.50$ $88.46$ $80.77$ $95.38$ $79.23$ $81.95$ $74.64$
o1-preview $93.08$ $95.38$ $97.69$ $96.15$ $86.92$ $86.15$ $90.00$ $63.08$ $72.50$ $85.00$ $88.46$ $89.23$ $98.46$ $96.15$ $89.59$ $87.31$
o1-mini $90.77$ $76.92$ $91.54$ $88.46$ $84.62$ $68.46$ $81.54$ $56.15$ $55.83$ $70.00$ $80.00$ $83.85$ $91.54$ $83.08$ $82.26$ $75.27$
zeroshot Input-Output prompt
LLAMA 405B $88.46$ $83.08$ $90.77$ $90.77$ $85.38$ $83.08$ $84.46$ $50.00$ $74.17$ $72.50$ $77.69$ $89.23$ $83.08$ $69.23$ $83.43$ $76.84$
GPT-4o Mini $70.77$ $66.92$ $68.46$ $80.77$ $80.00$ $58.46$ $54.62$ $21.54$ $57.50$ $55.83$ $56.92$ $44.62$ $64.62$ $66.15$ $64.7$ $56.33$
GPT-4o $68.46$ $83.08$ $71.54$ $84.62$ $74.62$ $77.69$ $56.15$ $37.69$ $60.00$ $69.17$ $59.23$ $86.92$ $76.92$ $80.00$ $66.7$ $74.17$
o1-preview $88.46$ $96.92$ $95.38$ $94.62$ $86.15$ $86.15$ $88.46$ $70.77$ $69.17$ $88.33$ $85.38$ $86.92$ $96.15$ $93.85$ $87.02$ $88.22$
o1-mini $90.77$ $96.92$ $93.08$ $92.31$ $83.85$ $79.23$ $83.08$ $54.62$ $55.00$ $60.83$ $90.00$ $89.23$ $95.38$ $91.54$ $84.45$ $80.67$
Table 6: Comparison of o1 Reasoning Model with the best performing LLMs on $7$ ACPBench tasks (Bool as well as MCQ). The right-most column is mean across tasks.
Recently, OpenAI released a series of LLM-based reasoning models called OpenAI o1 (OpenAI 2024b), that show significant improvement over GPT-4o on benchmarks that require reasoning. Although OpenAI o1 preview and mini are made available via similar APIs as previous LLMs, they do not truly fit the LLM category; rather, they are a system (or an agent) that makes multiple calls to LLMs before providing an answer. Note that OpenAI o1were also referred to as Large Reasoning Models (Valmeekam, Stechly, and Kambhampati 2024). While we acknowledge the difference, it is interesting to compare the best performing LLMs to the OpenAI o1 models. The comparison is depicted in Table 6. Further, Figure 7 shows the performance difference of OpenAI o1 models (with zeroshot IO and 2-shot COT prompts) from the best performing LLMs. Our results indicate that OpenAI o1 models fail to yield performance gains for boolean questions, but demonstrate notable improvements on MCQs. Specifically, OpenAI o1 preview consistently performs better or equal to the best performing model for MCQs. The responses for MCQ tasks suggests that OpenAI o1 models consider each option individually, perform a case-by-case analysis, and only then select an option.
We would like to reiterate that while we present results of OpenAI o1 side by side with LLAMA-3.1 405B and GPT-4o LLMs, the comparison is not even-handed due to below mentioned reasons:
  • All our LLM experiments had a generated token limit of $1024$; OpenAI o1 models did not have that limit. On average the number of tokens generated by OpenAI o1 preview for MCQ tasks, where we see the maximum improvement, was $5705$ (this includes the completion token ($3164$) and the reasoning token ($2542$)).
  • LLM evaluations are based on a single generation. We did not evaluate multi-turn prompts (such as self-consistency or self-reflection). OpenAI o1 models seem to internally make multiple calls to an LLM.
In terms of pure monetary cost, the OpenAI o1 evaluation is approximately $20$ times more expensive than of GPT-4o. It remains to be seen if a multi-turn prompting of an open-sourced LLM like LLAMA-3.1 can achieve similar improvement with lower cost.
@@ 5 Discussion and Future Work
In this work, we introduce ACPBench—a collection of datasets to evaluate the ability of LLMs to reason about action, change and planning. By evaluating $22$ state-of-the-art LLMs of varying size, we find these models underperform, even the largest ones, especially on tasks such as plan validation and action reachability. On the other hand, we show that finetuning a small language model, Granite 8B, can improve its reasoning ability to bring it on par with the best performing models. Further, we observe that the fine-tuned model exhibits remarkable generalization ability to unseen domains in ACPBench as well as to a different task in PlanBench. Further, our investigation with OpenAI o1 reasoning model indicates that OpenAI’s multi-turn approach yields improvements for multi-choice questions but fails to make an impact on boolean questions in ACPBench.
@@ A.2 Pretrained LLMs
Our paper presents domain-wise performance of few selected models in Figure  4, where we only presented results for the MCQ due to space constraints. In this section, we present the domain-wise analysis for all the $22$ pretrained LLMs for both boolean and multi-choice questions in Table 7.
In the main paper, Table 2 presents accuracy values with 2-shot COT prompting. We also attempted zeroshot Input Output (IO) prompt, the trends were similar. These zeroshot IO results aggregated over $5$ runs are presented in Table 8.
@@ A.3 Finetuned LLM
In Tables 3 and 4, we compare te finetuned model against the best performing model (last column). These values are obtained by looking at the aggregated performance on train and test domains respectively. We present performance of all the models on train and test domains in Tables 11 and  12.
@@ A.4 Ablation: Prompt Style
In our paper, we present an ablation to compare prompting styles. For that analysis; we presented results on test domains for multi-choice questions in Figure 5. Here, in Figure 8, we present performance on boolean questions.
@@
Figure 9: Comparison of different prompt styles on LLAMA 3.1 405B for Bool & MCQ tasks in all ACPBench domains.
@@
[ caption = Example questions for $7$ ACPBench tasks. For each question, an LLM was provided with context and then the question. The context contains natural language descriptions., label = tab:example_tasks, ] colspec = c—c—X—X, width = Task Context Bool Questions MCQ Questions
@@
Bool MCQ Bool MCQ Bool MCQ Bool MCQ Bool MCQ Bool MCQ Bool MCQ Bool MCQ Bool MCQ Bool MCQ Bool MCQ Bool MCQ Bool MCQ
@@
Table 7: Accuracy values for $22$ pretrained LLMs on Bool and MCQ tasks in ACPBench; segregated per domain.
@@
Table 8: Accuracy of $22$ LLMs on $7$ ACPBench tasks (Bool as well as MCQ) with zeroshot IO prompt. The right-most column is mean across tasks.